bad ←d 8%.de s.. (4) ent Pas-t ta isthe
Answer #1

a) WITHOUT REPLACEMENT:

Good=8

bad=4

total=12

total number of ways to select 3 eggs from total 12= 12C3

Possible ways of getting exactly two bad(out of 4) and one good(out of 8) = 4C2 * 8C1

7l n r!(-r) here, nCr-

Thus required probability= favorable outcomes/ total no. of outcomes

= (4C2 * 8C1) / 12C3

= 12!

=3 4 3*8 12 11 10

=0.21818

b) WITH REPLACEMENT:

total number of ways to select 3 eggs from total 12= 12^3

Possible ways of getting exactly two bad(out of 4) and one good(out of 8) :

GBB or BGB or BBG thus =

3 * 42 8

Thus required probability= favorable outcomes/ total no. of outcomes

=3 *42 81 12

0.638889